Suppose the production function is given by Q=L3\4K1\4 find APL and MPL function

1
Expert's answer
2022-02-07T08:33:41-0500

Solution:

APL = "\frac{Q}{L}"= "\frac{L^{3/4} K^{1/4} }{L} = \frac{K^{1/4} }{L^{1/4}}"

 

MPL = "\frac{\partial Q} {\partial L} = \frac{3} {4}L ^{-1/4}K^{1/4}"